Triose phosphate isomerase follows Michaelis-Menten kinetics. A 1.7 micromolar solution of triose phosphate isomerase has a kcat of 706 s-1 when converting glyceraldehyde-3-phosphate to dihydroxyacetone phosphate. What is triose phosphate isomerase's Vmax with glyceraldehyde-3-phosphate?

12 M-1 s-1
0.415 mmoles s-1
1.2 mmoles s-1
1.2 x 104 M-1 s-1
0.0024 moles s-1
Answer #1

Kcat= Vmax/ [ET]

[ET]= total enzyme concentration = 1.7 micromolar

Kcat= 706 s-1

[E] = 1.7 micromolar

Vmax= Kcat X [E]

Vmax= 706 s-1 X 1.7 micromolar = 1200.2 micromolar s-1

1 micromolar=0.001 mmol

Vmax= 1200.2 micromolar s-1 X 0.001 mmol= 1.2 mmol s-1 = 1.2 mmol/s

Right choice= 1.2 mmoles s-1